OpenStudy (anonymous):

ALGEBRA 1, PLEASE HELP. REWARDS WILL BE GIVEN. The graphs of functions f(x) and g(x) = f(x) + k are shown below: graph of line f of x going through ordered pairs 0, negative 2 and 2, 2. Graph of line g of x going through ordered pairs 0, 2 and 1, 4. The value of k is ___. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

Is the question complete? Just wondering since the blue asks box trails

OpenStudy (anonymous):

no that's the full question. @Sadworld

OpenStudy (anonymous):

I see now, give me a moment

OpenStudy (anonymous):

okay. take your time. :)

OpenStudy (anonymous):

f(x) = (x+k)+k f(x) = x+2k 0+2k = -2 (x, y of first order of pairs) k = -2/-2 k = -1

OpenStudy (anonymous):

thank you so much!@Sadworld

OpenStudy (anonymous):

wait, a min. Still proving

OpenStudy (anonymous):

You got an answer sheet?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So -1 is correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, lemme try another version.

OpenStudy (anonymous):

I got the substitutions wrong. f(x) = y so f(x) = f(x) + k -2 = -2 + k (y coordinate of the first pair) k = 1

OpenStudy (kittiwitti1):

@Directrix @ganeshie8

OpenStudy (anonymous):

hi Directrix, so I end up with k = 1. Wondering if I got it right.

Directrix (directrix):

No diagram?

OpenStudy (anonymous):



Directrix (directrix):

@Sadworld Did you get the equation for f(x) to be y = 2x -2 ?

OpenStudy (anonymous):

mmm... I got on my first try f(x) = x + 2k; On the 2nd try f(x) = y + k

Directrix (directrix):

And, the equation of g(x) = 2x + 2. From -2 to + 2 = 4 units for k but I need to make a graph because k may be -4.

OpenStudy (anonymous):

One part of the problem I encounter is the question. So [f(x) = f(x) + k] is wrong?

Directrix (directrix):

Look at the attached graphs from the Wolf. http://www.wolframalpha.com/input/?i=y+%3D+2x+%2B+2%2C+y+%3D+2x+-+2

OpenStudy (anonymous):

I knew it, I misunderstood the question.

Directrix (directrix):

I know what you are saying. That is why I asked to see the graph that came with the original problem. f(x) and g(x) = f(x) + k are the two graphs. This is not part of the problem: [f(x) = f(x) + k]

Directrix (directrix):

If the OP will post the original diagram, I'll reconsider. If not, I'm sticking with k = 4. It is as correct as I can figure. What do you think @Sadworld and what does @grunge101 think. Offline, it seems.

Directrix (directrix):

@kittiwitti1 Let us hear what you think, okay? All ideas are welcome and appreciated.

OpenStudy (anonymous):

I agree with you, I just jumbled to the question and thought that part was included to the whole thing.

OpenStudy (kittiwitti1):

I'm only on this to avoid getting repeated notifications about this question.

OpenStudy (kittiwitti1):

Lol I'm not inputting anything.

OpenStudy (anonymous):

lols

OpenStudy (anonymous):

OP will be so pissed to see all the additional chatter *laughs*

Directrix (directrix):

Look at my "wolf" graph. @kittiwitti1 The two functions are parallel and 4 apart?

Directrix (directrix):

I'm not saying that our OP did this but many OPs post a question and then run off.

OpenStudy (kittiwitti1):

Wolf graph?

OpenStudy (kittiwitti1):

Ah, that.

OpenStudy (kittiwitti1):

I am not participating in this. I am only here to minimize the notifications I get from this question lol

OpenStudy (kittiwitti1):

Sorry to be a bother. *bows*

Directrix (directrix):

I don't understand that @kittiwitti1

OpenStudy (anonymous):

You get computer messages if you participate in a question, participate in many you get overloaded by computer generated messages ie Spam

OpenStudy (kittiwitti1):

Pretty much what Sadworld said, yes.

Directrix (directrix):

Yes. But, I did not see the KW's name pinged in the thread. The OP did not ping me over here either.

OpenStudy (anonymous):

I asked Kitti for referrals since I only have 3 names i can refer

OpenStudy (kittiwitti1):

Yep. :p

Directrix (directrix):

Look at this. I found the solved problem AND the original diagram. http://openstudy.com/study#/updates/5398778de4b03871bbf0b21e

Directrix (directrix):

>>I asked Kitti for referrals since I only have 3 names i can refer I have seen people ping 12 people at the time. Why are you limited to three?

OpenStudy (anonymous):

My god, you actually spend time to dig it up o.o dedication.

OpenStudy (anonymous):

@Directrix I only ping people I know who are good ;)

Directrix (directrix):

Well, that is flattering.

Directrix (directrix):

We can close this up now with a big FOUR, can't we?

Directrix (directrix):

Correct.

OpenStudy (anonymous):

The answer is \[k = 4\]

Directrix (directrix):

Let's go here and help. http://openstudy.com/study#/updates/54a2debee4b0a96030d794d1